# Bayes Adaptive Monte Carlo Tree Search for Offline Model-based Reinforcement Learning

### v1

@@ Appendix C Key Hyperparameter Setup
$K$ $\lambda$ $H$ $N$ $K$ $\lambda$ $H$ $N$ $K$ $\lambda$ $H$ $N$
@@
Table 3: Key hyperparameters of the proposed algorithms for each evaluation task. $K$: ensemble size, $\lambda$: reward penalty coefficient, $H$: rollout horizon, $N$: number of training epochs.
In Table 3, we list the key hyperparameters of the proposed algorithms. For each task, an ensemble of $K$ dynamics and reward models is trained using the provided offline dataset. These learned models are then utilized as a simulator to train a control policy using off-the-shelf RL methods, such as SAC. The policy is trained for $N$ epochs. At each epoch, $50000H$ transitions are sampled by interacting with the simulator, followed by $1000$ RL training iterations. In particular, $50000$ states are randomly sampled from the offline dataset, with each state followed by a rollout lasting $H$ time steps. To mitigate overestimation, a reward penalty based on the discrepancy among the ensemble members is applied with a coefficient $\lambda$, as shown in Equation (5). The setups for $K$, $\lambda$, and $H$ are almost the same across the three algorithms and primarily inherited from the baseline – “Optimized" (Lu et al. (2022)), to make sure the improvements are brought by the Bayesian RL and deep search components.
The policy is evaluated on the ground truth environment for 10 episodes at the end of each training epoch. We report the average scores across the final 10 training epochs of our algorithms in Tables 1 and 5. It is important to note that increasing the number of training epochs $N$ does not necessarily lead to better policy performance, since the training is based on learned dynamics and reward models rather than the ground truth. According to (Lu et al. (2022)) and our experiments, the hyperparameters listed above can significantly influence the performance of model-based RL. Adjusting these hyperparameters could either enhance or impair the learning performance of our algorithms. We also suspect that the performance of the baselines listed in Tables 1 and 5, which are from their original papers, could be further improved by fine-tuning the relevant hyperparameters.
@@
Figure 3: Performance of our proposed algorithms on D4RL MuJoCo tasks. The results for HalfCheetah, Hopper, and Walker2d are presented in the three rows, respectively. Each subfigure depicts the change in the undiscounted episodic return as a function of training epochs. Experiments are repeated three times with different random seeds, with the solid line representing the mean and the shaded area indicating the 95% confidence interval. For reference, the expert-level episodic returns for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively. Note that the training epochs for each algorithm, as listed in Table 3, have been linearly scaled to 800 for better visualization.
BA-MCTS and BA-MCTS-SL utilize Bayes Adaptive Monte Carlo Tree Search to collect samples for offline model-based RL. Instead of performing a tree search at every state, we randomly select a proportion (i.e., $\rho$) of states from the available 50000 states at each rollout time step as root nodes for tree search. For the remaining states, actions are sampled directly from the policy, i.e., $a\sim\pi(\cdot|s)$. The tree search procedure is detailed in Algorithm 1, with the number of MCTS iterations, $E$, set to 50. Increasing $\rho$ and $E$ can potentially enhance performance, but it will also linearly increase the computational cost. Table 4 outlines the key hyperparameters related to the search process for each algorithm and task. (1) As described in Algorithm 1, the parameters $\alpha$ and $\beta$ control the rate of double progressive widening. ($\beta$ is set as 0.5 across all tasks.) To encourage deeper search, we limit the number of actions sampled from a state under $n_{a}$ and the number of next states sampled from an action under $n_{s}$, respectively. Action selection follows the UCT rule, as discussed in Appendix A, where $c>0$ balances the exploration and exploitation. Additionally, inspired by the success of MuZero in enhancing exploration, we introduce Dirichlet noise $x_{d}$ at the root nodes, where actions are sampled from a mixture of distributions: $a\sim\eta x_{d}+(1-\eta)\pi(\cdot|s)$ and $\eta$ controls the mixture rate. Notably, for $(c,\eta,n_{a},n_{s})$, we explore the set of possible combinations: $\{(2.5,0.3,20,1),(1.0,0.3,20,1),(1.0,0.1,20,1),(1.0,0.3,10,5),(1.0,0.1,10,5)\}$ during hyperparameter fine-tuning. We believe there are likely more optimal search settings yet to be discovered. (2) In BA-MCTS-SL, policy improvement is achieved through supervised learning. We find that, rather than learning solely from samples collected within the current epoch, incorporating a buffer of samples from the past $N_{SL}$ epochs helps to stabilize the learning process. Also, in the Walker2d environment, BA-MCTS-SL requires a warm-up training phase of $N_{P}$ epochs using BA-MBRL, allowing the initial policy to generate effective signals for supervised learning.
Appendix D Comparisons with Model-free Methods on D4RL MuJoCo
@@
Data Type Environment BA-MCTS -SL (ours) BA-MCTS (ours) BA-MBRL (ours) CQL BEAR BRAC-v SAC BC
random HalfCheetah 29.20 $\pm$ 2.00 36.23 $\pm$ 1.04 32.76 $\pm$ 1.16 35.4 25.1 31.2 30.5 2.1
random Hopper 33.83 $\pm$ 0.10 31.56 $\pm$ 0.12 31.47 $\pm$ 0.03 10.8 11.4 12.2 11.3 1.6
random Walker2d 21.89 $\pm$ 0.07 21.59 $\pm$ 0.32 21.45 $\pm$ 0.53 7.0 7.3 1.9 4.1 9.8
medium HalfCheetah 70.47 $\pm$ 3.52 75.84 $\pm$ 3.81 56.54 $\pm$ 5.20 44.4 41.7 46.3 -4.3 36.1
medium Hopper 97.75 $\pm$ 7.09 96.70 $\pm$ 14.0 98.25 $\pm$ 3.42 86.6 52.1 31.1 0.8 29.0
medium Walker2d 82.24 $\pm$ 1.85 74.73 $\pm$ 3.25 75.41 $\pm$ 4.17 74.5 59.1 81.1 0.9 6.6
med-replay HalfCheetah 61.16 $\pm$ 1.60 65.45 $\pm$ 0.81 62.50 $\pm$ 0.18 46.2 38.6 47.7 -2.4 38.4
med-replay Hopper 106.3 $\pm$ 0.13 101.8 $\pm$ 3.46 93.91 $\pm$ 4.25 48.6 33.7 0.6 3.5 11.8
med-replay Walker2d 92.13 $\pm$ 5.13 95.06 $\pm$ 2.11 97.54$\pm$ 1.93 32.6 19.2 0.9 1.9 11.3
med-expert HalfCheetah 80.53 $\pm$ 6.63 76.16 $\pm$ 10.3 90.52 $\pm$ 4.13 62.4 53.4 41.9 1.8 35.8
med-expert Hopper 112.2 $\pm$ 0.29 108.3 $\pm$ 0.22 107.8 $\pm$ 0.37 111 96.3 0.8 1.6 111.9
med-expert Walker2d 107.7 $\pm$ 0.82 110.0 $\pm$ 1.74 84.71 $\pm$ 0.87 98.7 40.1 81.6 -0.1 6.4
Average Score 74.62 74.45 71.06 54.85 39.83 31.44 4.13 25.07
Table 5: Comparisons between the proposed algorithms and model-free offline policy learning methods on the D4RL benchmark suite. Each value represents the normalized score, as proposed in (Fu et al. (2020)), of the policy trained by the corresponding algorithm. These scores are undiscounted returns normalized to approximately range between 0 and 100, where a score of 0 corresponds to a random policy and a score of 100 corresponds to an expert-level policy. For our algorithms, we report the average score of the final ten policy learning epochs and its standard deviation across three different random seeds.
As a complement to Table 1, we compare our algorithms with a series of model-free offline policy learning (Chen et al. (2024)) methods. We include SOTA model-free offline RL methods: CQL (Kumar et al. (2020)), BEAR (Kumar et al. (2019)), and BRAC-v (Wu et al. (2019)). Additionally, we show the performance of directly applying SAC or Behavioral Cloning (BC, Chen et al. (2024)) to the provided offline dataset in the last two columns. The mean performance of the baselines are taken from related works (Yu et al. (2020); Kidambi et al. (2020); Fu et al. (2020)). Our algorithms show significantly better performance, demonstrating the necessity of model-based learning in these environments. The training plots of our proposed algorithms in each environment is further detailed in Figure 3.
Appendix E Computation Cost of Sampled EfficientZero on D4RL MuJoCo
Here, we report the training time of Sampled EfficientZero on the D4RL MuJoCo tasks in Table 6. This set of experiments is run on a server with 40 Intel(R) Xeon(R) Gold 5215 CPUs and 4 Tesla V100-SXM2-32GB GPUs. Despite the substantial computational input, no corresponding performance improvement is observed. We welcome researchers to use our released code for further hyperparameter fine-tuning and to improve the results we have provided.
hc-med-expert hc-med-replay hc-medium hc-random
55.4 $\pm$ 1.3 54.8 $\pm$ 1.8 54.7 $\pm$ 1.9 54.8 $\pm$ 2.6
hp-med-expert hp-med-replay hp-medium hp-random
66.1 $\pm$ 0.9 126.6 $\pm$ 12.0 70.5 $\pm$ 1.0 153.7 $\pm$ 19.4
wk-med-expert wk-med-replay wk-medium wk-random
60.8 $\pm$ 1.8 134.9 $\pm$ 4.2 63.7 $\pm$ 1.5 175.7 $\pm$ 28.2

### v4

@@ Appendix D Key Hyperparameter Setup
$K$ $\lambda$ $H$ NN $K$ $\lambda$ $H$ NN $K$ $\lambda$ $H$ NN
@@
Table 5: Key hyperparameters of the proposed algorithms for each evaluation task in the ablation study. $K$: ensemble size, $\lambda$: reward penalty coefficient, $H$: rollout horizon, NN: number of training epochs.
Table 5 lists the key hyperparameters of BA-MCTS and its ablated variants used in the ablation study. For each task, an ensemble of $K$ dynamics and reward models is trained using the provided offline dataset. These learned models are then utilized as a simulator to train a control policy using off-the-shelf RL methods, such as SAC. The policy is trained for $N$ epochs. At each epoch, $50000H$ transitions are sampled by interacting with the simulator, followed by $1000$ RL training iterations. In particular, $50000$ states are randomly sampled from the offline dataset, with each state followed by a rollout lasting $H$ time steps. To mitigate overestimation, a reward penalty based on the discrepancy among the ensemble members is applied with a coefficient $\lambda$, as shown in Equation (5). The setups for $K$, $\lambda$, and $H$ are almost the same across the three algorithms and primarily inherited from the baseline – “Optimized” (Lu et al., 2022), to make sure the improvements are brought by the Bayesian RL and deep search components (rather than tricky hyperparameter setups).
The policy is evaluated on the ground truth environment for 10 episodes at the end of each training epoch. For all tables involving benchmarking results on D4RL, we report the average scores across the final 10 training epochs of our algorithms. It is important to note that increasing the number of training epochs $N$ does not necessarily lead to better policy performance, since the training is based on learned dynamics and reward models rather than the ground truth. According to Lu et al. (2022) and our experiments, the hyperparameters listed above can significantly influence the performance of model-based RL. Adjusting these hyperparameters could either enhance or impair the learning performance of our algorithms. We also suspect that the performance of the baselines listed in Table 9, which are from their original papers, could be further improved by fine-tuning the relevant hyperparameters.
@@
BA-MCTS and BA-MCTS-SL utilize Continuous BAMCP (i.e., Algorithm 1) to collect samples for subsequent policy distillation. Instead of performing a tree search at every state, we randomly select a proportion (i.e., $\rho$) of states from the available 50000 states at each rollout time step as root nodes for tree search to decide on the optimal actions. For the remaining states, actions are sampled directly from the policy, i.e., $a\sim\pi(\cdot|s)$. The tree search procedure is detailed in Algorithm 1, with the number of MCTS iterations, $E$, set to 50. Increasing $\rho$ and $E$ can potentially enhance performance, but it will also linearly increase the computational cost. Table 6 outlines the key hyperparameters related to the search process for each algorithm and task. (1) As described in Algorithm 1, the parameters $\alpha$ and $\beta$ control the rate of double progressive widening. ($\beta$ is set as 0.5 across all tasks.) To encourage deeper search, we limit the number of actions sampled from a state under $n_{a}$ and the number of next states sampled from an action under $n_{s}$, respectively. Action selection follows the UCT rule, as discussed in Appendix A, where $c>0$ balances the exploration and exploitation. Additionally, inspired by the success of MuZero in enhancing exploration, we introduce Dirichlet noise $x_{d}$ at the root nodes, where actions are sampled from a mixture of distributions: $a\sim\eta x_{d}+(1-\eta)\pi(\cdot|s)$ and $\eta$ controls the mixture rate. Notably, for $(c,\eta,n_{a},n_{s})$, we explore the set of possible combinations: $\{(2.5,0.3,20,1),(1.0,0.3,20,1),(1.0,0.1,20,1),(1.0,0.3,10,5),(1.0,0.1,10,5)\}$ during hyperparameter fine-tuning. We believe there are likely more optimal search settings yet to be discovered. (2) In BA-MCTS-SL, policy improvement is achieved through supervised learning. We find that, rather than learning solely from samples collected within the current epoch, incorporating a buffer of samples from the past $N_{SL}$ epochs helps to stabilize the learning process. Also, in the Walker2d environment, BA-MCTS-SL requires a warm-up training phase of $N_{P}$ epochs using BA-MBRL, allowing the initial policy to generate effective signals for supervised learning.
Appendix E Details of the Tokamak Control Tasks
STATE SPACE
ACTION SPACE
Targets Current Target, Density Target
Table 7: The state and action spaces of the tokamak control tasks.
Nuclear fusion is a promising energy source to meet the world’s growing demand. It involves fusing the nuclei of two light atoms, such as hydrogen, to form a heavier nucleus, typically helium, releasing energy in the process. The primary challenge of fusion is confining a plasma, i.e., an ionized gas of hydrogen isotopes, while heating it and increasing its pressure to initiate and sustain fusion reactions. The tokamak is one of the most promising confinement devices. It uses magnetic fields acting on hydrogen atoms that have been ionized (given a charge) so that the magnetic fields can exert a force on the moving particles (Pironti and Walker, 2005).
Char et al. (2024) trained a deep recurrent network as a dynamics model for the DIII-D tokamak, a device located in San Diego, California, and operated by General Atomics, using a large dataset of operational data from that device. A typical shot (i.e., episode) on DIII-D lasts around 6-8 seconds, consisting of a one-second ramp-up phase, a multi-second flat-top phase, and a one-second ramp-down phase. The DIII-D also features several real-time and post-shot diagnostics that measure the magnetic equilibrium and plasma parameters with high temporal resolution. The authors demonstrate that the learned model predicts these measurements for entire shots with remarkable accuracy. Thus, we use this model as a “ground truth” simulator for tokamak control tasks. Specifically, we generate a dataset of 725270 transitions for offline RL and evaluate the learned policy using this data-driven simulator.
The state and action spaces for the tokamak control tasks are outlined in Table 7. For detailed physical explanations of their components, please refer to Abbate et al. (2021); Char et al. (2023); Ariola et al. (2008). The state space consists of five scalar values and six profiles which are discretized measurements of physical quantities along the minor radius of the toroid. After applying principal component analysis (Maćkiewicz and Ratajczak, 1993), the pressure profile is reduced to two dimensions, while the other profiles are reduced to four dimensions each. In total, the state space comprises 27 dimensions. The action space includes direct control actuators for neutral beam power, torque, gas, ECH power, current, and magnetic field, as well as target values for plasma density and plasma shape, which are managed through a lower-level control module. Altogether, the action space consists of 14 dimensions. While for certain tasks, it is possible to prune the state and action spaces to reduce the learning complexity, we have chosen not to apply any domain-specific knowledge in these evaluations for general RL algorithms. We reserve the domain-specific applications of our algorithms, which would require more domain knowledge and engineering efforts, as an important future work.
We select a reference shot from DIII-D, which spans 251 time steps, and use its trajectories of Ion Rotation, Electron Temperature, and $\beta_{n}$ as targets for three tracking tasks. Specifically, $\beta_{n}$ is the normalized ratio between plasma pressure and magnetic pressure, a key quantity serving as a rough economic indicator of efficiency. Since the tracking targets vary over time, we include the time step as part of the policy input. The reward function for each task is defined as the negative squared tracking error of the corresponding component (i.e., temperature, rotation, or $\beta_{n}$) at each time step, and the reward is normalized by the episode horizon (i.e., 251 time steps). Notably, for policy learning, the reward function is provided rather than learned from the offline dataset as in D4RL tasks; and the dataset (for offline RL) does not include the reference shot or any nearby, similar shots, making pure imitation infeasible.
Appendix F Additional Evaluation Results
This appendix provides detailed empirical results and analyses to supplement the findings presented in Section 5. Appendix F.1 offers both quantitative and qualitative analyses of the benefits of belief adaptation, presenting additional results for Section 5.1.
The following appendices report additional experimental results for benchmarking on MuJoCo D4RL tasks (i.e., Section 5.2):
  • Appendix F.2: Evaluates our algorithm against a suite of model-free offline RL methods.
  • Appendix F.3: Presents a comprehensive comparison against 19 offline RL baselines, including a statistical significance analysis.
  • Appendix F.4: Provides a detailed comparison with MuZero and its variants, focusing on their algorithm design, empirical performance, and computational cost.
Finally, we provide complete results for the ablation study (i.e., Section 5.3):
  • Appendix F.5: Demonstrates the significant performance enhancement achieved by the belief adaptation module and our MCTS-based planner.
  • Appendix F.6: Discusses the trade-offs between the two policy distillation mechanisms investigated (i.e., SAC and supervised learning).
  • Appendix F.7: Presents an ablation study on the necessity of using reward penalty in offline MBRL.
F.1 Quantitative and Qualitative Analysis of the Effectiveness of Belief Adaptation
To understand why belief adaptation over an ensemble of models using a Bayes-Adaptive MDP (BAMDP) is effective, we analyze its impact on the prediction accuracy of the learned world models. The Bayesian adaptation, as defined in Eq. (4), uses observed state transition sequences to adjust the belief over each ensemble member, thereby enhancing the quality of predicted rollout trajectories.
For each benchmarking task in D4RL MuJoCo, we compute the average transition likelihood (i.e., $\mathcal{P}(s^{\prime}|s,a)\mathcal{R}(r|s,a)$) for the provided offline rollouts, comparing cases with and without belief adaptation. The ratios of these likelihoods are presented in Table 8. Results show that the offline rollouts, collected from real environments, are more likely under the adapted ensemble. This demonstrates that Bayesian belief adaptation serves as an effective calibration method for the learned world models.
hc-med-expert hc-med-replay hc-medium hc-random
0.7515 2.6315 2.2674 2.4561
hp-med-expert hp-med-replay hp-medium hp-random
14.064 4.2214 3.5913 1.7400
wk-med-expert wk-med-replay wk-medium wk-random
2.0898 3.8169 34.264 1.0341
Table 8: Ratios of average transition likelihoods in offline data with and without belief adaptation across ensemble members. A ratio greater than 1 indicates that the adapted ensemble assigns higher likelihood to the true transitions.
Furthermore, we evaluate the quality of imaginary rollouts generated by the learned models. Table 1 in the main content presents the prediction errors for next states and rewards in these rollouts. The results reveal that the adaptive ensemble achieves more accurate predictions in 10 out of 12 benchmarking tasks, indicating that belief adaptation leads to a more accurate model for planning.
(a) Offline rollout
(b) Imaginary rollout (at step 0)
(c) Imaginary rollout (at step 22)
(d) Imaginary rollout (at step 495)
Figure 1: Belief adaptation during offline and imaginary rollouts. (a) shows the belief over twelve ensemble members, adapting to an offline trajectory of Hopper-med-expert. (b), (c), and (d) illustrate the belief changes during imaginary rollouts which start from different points of the offline trajectory shown in (a).
As a qualitative analysis, we plot the belief adaptation over an offline trajectory (from $\mathcal{D}_{\mu}$) in Hopper-med-expert in Figure 1(a). Initially, all twelve ensemble members have the same belief. As the trajectory progresses, the beliefs of each member are updated based on the transition history, with the dominant model (the one with the highest belief) continuously changing. Figure 1(b)-1(d) further show the belief adaption during imaginary rollouts that start from different time steps of that offline trajectory.
F.2 Comparison with Model-free Methods on D4RL MuJoCo
Data Type Environment BA-MCTS (ours) CQL BEAR BRAC-v SAC BC
random HalfCheetah 41.45 (1.40) 35.4 25.1 31.2 30.5 2.1
random Hopper 31.42 (0.10) 10.8 11.4 12.2 11.3 1.6
random Walker2d 21.54 (0.06) 7.0 7.3 1.9 4.1 9.8
medium HalfCheetah 77.31 (1.40) 44.4 41.7 46.3 -4.3 36.1
medium Hopper 103.9 (0.33) 86.6 52.1 31.1 0.8 29.0
medium Walker2d 88.23 (0.78) 74.5 59.1 81.1 0.9 6.6
med-replay HalfCheetah 71.24 (2.65) 46.2 38.6 47.7 -2.4 38.4
med-replay Hopper 106.4 (0.53) 48.6 33.7 0.6 3.5 11.8
med-replay Walker2d 91.42 (1.54) 32.6 19.2 0.9 1.9 11.3
med-expert HalfCheetah 102.3 (2.27) 62.4 53.4 41.9 1.8 35.8
med-expert Hopper 111.8 (0.82) 111 96.3 0.8 1.6 111.9
med-expert Walker2d 116.0 (1.49) 98.7 40.1 81.6 -0.1 6.4
Average Score 80.25 54.85 39.83 31.44 4.13 25.07
Table 9: Comparisons between the proposed algorithms and model-free offline policy learning methods on the D4RL benchmark suite. Each value represents the normalized score, as proposed in Fu et al. (2020), of the policy trained by the corresponding algorithm. These scores are undiscounted returns normalized to approximately range between 0 and 100, where a score of 0 corresponds to a random policy and a score of 100 corresponds to an expert-level policy. For our algorithms, we report the average score of the final ten policy learning epochs and its standard deviation across three different random seeds.
Here, we compare our algorithms with a series of model-free offline policy learning (Chen et al., 2024) methods in Table 9. We include SOTA model-free offline RL methods: CQL (Kumar et al., 2020), BEAR Kumar et al. (2019), and BRAC-v (Wu et al., 2019). Additionally, we show the performance of directly applying SAC or Behavioral Cloning (BC) (Chen et al., 2024) to the provided offline dataset in the last two columns. The mean performance of the baselines are taken from related works (Yu et al., 2020; Kidambi et al., 2020; Fu et al., 2020). Our algorithm shows significantly better performance, demonstrating the necessity of model-based learning in these environments.
@@ F.2 Comparison with Model-free Methods on D4RL MuJoCo
Figure 2: Performance of Sampled EfficientZero on D4RL MuJoCo tasks. The results for HalfCheetah, Hopper, and Walker2d are presented in the three rows, respectively. Each subfigure depicts the change in undiscounted episodic return as a function of the number of training samples. Experiments are repeated three times with different random seeds, with the solid line representing the mean and the shaded area indicating the 95% confidence interval. For reference, the expert-level episodic returns (corresponding to scores of 100) for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively.
F.3 Comprehensive Benchmarking on D4RL MuJoCo
To firmly establish the state-of-the-art (SOTA) performance of our algorithm, this section provides a comprehensive comparison of BA-MCTS against 19 offline RL methods on the 12 D4RL MuJoCo tasks. To provide context, we briefly categorize the baselines:
  • The model-based (MBRL) methods include classic approaches like MOPO (Yu et al., 2020), MOReL (Kidambi et al., 2020), and COMBO (Yu et al., 2021), as well as more recent/SOTA methods such as MOBILE (Sun et al., 2023), CBOP (Jeong et al., 2023), RAMBO (Rigter et al., 2022), and MAPLE (Chen et al., 2021b).
  • The model-free (MFRL) methods cover a wide range of techniques. We include prominent SOTA methods such as the value-based algorithm CQL (Kumar et al., 2020), the implicit Q-learning method IQL (Kostrikov et al., 2021), ensemble-based methods like SAC-N and EDAC (An et al., 2021), the policy constraint method TD3+BC (Fujimoto and Gu, 2021), and the sequence modeling approach Decision Transformer (DT) (Chen et al., 2021a). We also compare against classic MFRL baselines, including BEAR (Kumar et al., 2019), BRAC-v (Wu et al., 2019), and the direct application of SAC and Behavioral Cloning (BC) (Chen et al., 2024). Finally, APE-V (Ghosh et al., 2022) is included as a conceptually related model-free method that also leverages the BAMDP framework.
The results, including a statistical significance analysis using Cohen’s d (Cohen, 1988), are summarized in Table 3 in the main content. Note that MOReL, IQL, and the model-free offline RL baselines in Table 9 did not report standard deviations. Additionally, IQL and Decision Transformer (DT) were not evaluated on D4RL tasks with random data.
As shown in Table 3, BA-MCTS achieves the highest average score among all 20 algorithms, both including and excluding the challenging ‘random‘ datasets. To quantify the significance of this improvement, we compute Cohen’s d against all baselines for which standard deviations are available. The results show that the Cohen’s d value is greater than 1.8 in all cases, and for most baselines, it is significantly larger (e.g., >> 3.5). This indicates that the performance improvement of our algorithm is not only substantial in magnitude but also statistically significant. This comprehensive comparison provides strong evidence that our proposed method establishes a new state-of-the-art in offline RL on this widely-used benchmark.
F.4 Comparison with MuZero and its Variants
Monte-Carlo Tree Search (MCTS) (Browne et al., 2012) has been successfully integrated with RL, as exemplified by AlphaZero (Silver et al., 2017) and MuZero (Schrittwieser et al., 2020). These methods have achieved superhuman performance in domains requiring highly sophisticated decision-making processes. AlphaZero relies on a given world model, whereas MuZero learns the world model and policy simultaneously by interacting with the environment. Although there have been various extensions of MuZero (Hubert et al., 2021; Schrittwieser et al., 2021; Ye et al., 2021; Antonoglou et al., 2022; Oren et al., 2022; Zhao et al., 2024), most are designed for online MBRL. According to Niu et al. (2023), applying MuZero to offline learning, especially for continuous control in highly stochastic environments, remains a significant open challenge.
Our algorithm design differs from MuZero (and its variants) in several key aspects: (1) MuZero integrates model learning and policy training into a single stage, using a world model defined in a latent state space. Our algorithm separately learns a world model and then trains the policy on top of it, aligning with the widely-adopted offline MBRL framework. (2) MuZero employs a single latent world model (rather than an ensemble) and does not explicitly account for uncertainty in dynamics or reward predictions. This makes it particularly vulnerable to modeling errors in offline MBRL. (3) We introduce double progressive widening (Auger et al., 2013) and Bayes-adaptive planning into MCTS, resulting in a fundamentally different planning algorithm compared to the one used in MuZero.
To evaluate the performance of MuZero-type methods on D4RL MuJoCo tasks, we use the open-source implementation and hyperparameter configurations of Sampled EfficientZero (Ye et al., 2021) provided by LightZero (Niu et al., 2023). Benchmarking results from LightZero indicate that Sampled EfficientZero achieves the best performance on (online) MuJoCo locomotion tasks compared to other MuZero variants. To adapt it for offline learning, we employ the reanalyse technique proposed by Schrittwieser et al. (2021).
The evaluation results are presented in Figure 2. As shown, the performance is significantly worse compared to our methods listed in Figure 3, despite Sampled EfficientZero’s higher computational cost. For reference, the expert-level episodic returns (corresponding to scores of 100) for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively.
In Table 10, we report the training time of our proposed algorithms and Sampled EfficientZero on the D4RL MuJoCo tasks. The experiments were conducted on a server with 40 Intel(R) Xeon(R) Gold 5215 CPUs and 4 Tesla V100-SXM2-32GB GPUs. Our search-based algorithm, BA-MCTS, requires considerably less training time than Sampled EfficientZero to achieve its superior performance. While BA-MCTS requires more computation than the non-search variant BA-MBRL, this extra cost is limited to the offline training phase; no MCTS is performed during deployment, ensuring real-time execution.
@@ F.4 Comparison with MuZero and its Variants
Data Type Environment BA-MCTS -SL (ours) BA-MCTS (ours) BA-MBRL (ours) Sampled EfficientZero
random HalfCheetah 11.2 (2.6) 14.6 (2.2) 1.2 (0.4) 54.8 (2.6)
random Hopper 48.7 (2.4) 65.2 (1.9) 5.6 (0.8) 153.7 (19.4)
random Walker2d 25.7 (2.0) 38.1 (1.1) 5.2 (1.1) 175.7 (28.2)
medium HalfCheetah 12.6 (3.0) 15.3 (1.3) 1.9 (0.2) 54.7 (1.9)
medium Hopper 18.9 (4.8) 67.5 (0.2) 1.8 (0.2) 70.5 (1.0)
medium Walker2d 24.0 (1.8) 33.3 (4.0) 4.8 (1.0) 63.7 (1.5)
med-replay HalfCheetah 24.7 (0.9) 22.4 (1.2) 2.1 (0.4) 54.8 (1.8)
med-replay Hopper 17.9 (0.0) 12.7 (1.2) 5.4 (0.2) 126.6 (12.0)
med-replay Walker2d 40.3 (7.2) 77.0 (2.4) 8.0 (0.3) 134.9 (4.2)
med-expert HalfCheetah 32.6 (0.1) 12.6 (0.8) 4.7 (1.1) 55.4 (1.3)
med-expert Hopper 61.5 (4.6) 76.9 (11.6) 5.3 (0.5) 66.1 (0.9)
med-expert Walker2d 35.0 (6.5) 55.9 (1.8) 2.9 (0.9) 60.8 (1.8)